OpenStudy (anonymous):

Solve the proportion for x: 8x/10-2=5x+3/7x Can someone please explain?

OpenStudy (anonymous):

how?

jimthompson5910 (jim_thompson5910):

if you have A/B = C/D then you can cross multiply to get A*D = B*C

jimthompson5910 (jim_thompson5910):